# Geometric and Dynamic Scaling in Deep Transformers

### v2

@@ Abstract
Scaling Transformer architectures to extreme depth often leads to rank collapse: representations become redundant and degenerate despite modern normalization schemes. We argue this is fundamentally a geometric problem. Standard residual connections implicitly assume monotonic feature accumulation is beneficial, but provide no mechanism to constrain update directions or erase outdated information. As depth increases, this causes uncontrolled drift from the semantic manifold and representational collapse.
We propose the Manifold-Geometric Transformer (MGT), a unified framework addressing these failures through two orthogonal principles. First, manifold-constrained hyper-connections (mHC) restrict residual updates to valid tangent space directions, preventing manifold drift. Second, deep delta learning (DDL) enables data-dependent, non-monotonic updates that support feature erasure rather than unconditional accumulation. Together, mHC controls update direction while DDL controls magnitude and sign, yielding stable geometric evolution across depth.
@@ 1 Introduction
Simultaneously, the additive nature implies a “write-only” memory mechanism. It is optimizationally difficult for a network to “erase” information (i.e., learn $\mathcal{F}(\mathbf{x})\approx-\mathbf{x}$) when context shifts. This results in the “residual accumulation” problem, where noise accumulates over depth.
@@ 2 Related Work
Rank Collapse in Deep Networks. dong2021attention provided seminal theoretical analysis showing that pure self-attention layers lose rank doubly exponentially with depth, fundamentally limiting representational capacity. This collapse occurs even with modern normalization techniques (ba2016layer; xiong2020layer), suggesting that architectural modifications beyond optimization tricks are necessary. Recent depth-scaling efforts (wang2022deepnet; Touvron_2021_ICCV) have proposed specialized initialization and normalization schemes to train networks exceeding 100 layers, yet these approaches treat symptoms rather than addressing the geometric root cause of collapse.
Residual Connections and Depth. Since ResNets (he2016deep), residual connections have enabled training of deep networks by providing gradient highways. However, the standard formulation $\mathbf{x}_{l+1}=\mathbf{x}_{l}+\mathcal{F}(\mathbf{x}_{l})$ implicitly assumes additive updates in Euclidean space are universally beneficial. As Touvron_2021_ICCV observed, this monotonic accumulation can lead to feature redundancy in very deep networks. Our work addresses this through explicit erasure mechanisms.
Manifold-Constrained Hyper-Connections (mHC). The recently proposed mHC framework (xie2025mhc) addresses stability in width-scaling by constraining hyper-connections to geometric manifolds (e.g., the Birkhoff polytope), preserving the identity mapping property of ResNets. We adopt this perspective as a “spatial regularizer” that projects updates onto the tangent space $T_{\mathbf{x}}\mathcal{M}$ of the data manifold. While xie2025mhc focused on connection geometry, we argue that dynamics along these connections require complementary mechanisms for feature redundancy.
Deep Delta Learning (DDL). Our dynamic update law builds on the recent Deep Delta Learning framework (zhang2026deep), which generalizes residual connections via a learnable Delta Operator—a rank-1 perturbation parameterized by reflection direction $k(\mathbf{x})$ and gate $\beta(\mathbf{x})$. The key insight is that $\beta$ enables interpolation between identity ($\beta\to 0$), projection ($\beta\to 1$), and reflection ($\beta\to 2$), providing an explicit mechanism for information erasure that allows networks to actively forget outdated features.
@@ 3.1 Architecture Overview
We redesign the fundamental building block of the Transformer to explicitly separate feature generation from feature propagation. As illustrated in Figure 1, the proposed Manifold-Geometric Transformer (MGT) Block deviates from the standard Post-LN structure by introducing a geometric processing stage before the residual addition.
@@
[…]
Figure 1: Architecture of the Manifold-Geometric Transformer (MGT) Block. The pipeline explicitly separates (1) feature generation, (2) geometric rectification (blue/purple), and (3) dynamic erasure (orange). The dashed orange line illustrates the context-aware gating mechanism derived directly from the input state.
@@ 3.2 Preliminaries: The Manifold Hypothesis in Deep Layers
Let $\mathcal{X}=\{\mathbf{x}_{l}\}_{l=0}^{L}$ denote the sequence of hidden states in a Transformer, where $\mathbf{x}_{l}\in\mathbb{R}^{S\times D}$. We posit that valid semantic representations do not populate the entire Euclidean space $\mathbb{R}^{D}$, but rather lie on a lower-dimensional Riemannian manifold $\mathcal{M}\subset\mathbb{R}^{D}$. A naive step $\mathbf{x}_{l}+\mathbf{v}_{l}$ often forces the representation off the manifold ($\mathbf{x}_{l+1}\notin\mathcal{M}$). To maintain structural integrity, the update vector must be constrained to the local tangent space $T_{\mathbf{x}_{l}}\mathcal{M}$.
@@ 3.3 Manifold-Constrained Hyper-Connections (mHC)
The goal of mHC is to rectify the raw update vector $\mathbf{v}_{raw}$ generated by the sub-layers (MHSA or FFN). Since computing the exact tangent space via eigendecomposition is computationally prohibitive, we employ an efficient Soft Subspace Approximation.
@@
$\mathbf{v}_{mHC}=\mathbf{v}_{raw}\odot\sigma(\text{LN}(\mathbf{W}_{gate}\mathbf{x}_{l}))$ (1)
where $\mathbf{W}_{gate}$ projects the input to a semantic importance score. While not a strict orthogonal projection in the linear algebra sense, this acts as a soft manifold constraint, suppressing feature dimensions (noise subspace) that deviate from the current semantic trajectory defined by $\mathbf{x}_{l}$.
3.4 Deep Delta Learning (DDL): The Geometric Update Law
Once the update direction is rectified by mHC, we must determine the update dynamics. Standard connections assume a fixed step size. To enable true erasure (moving backwards along the current state vector), we adopt the generalized Householder form (zhang2026deep).
We introduce the Delta Controller $\beta(\mathbf{x}_{l})$:
$\bm{\beta}=\lambda\cdot\tanh(\mathcal{G}(\mathbf{x}_{l}))+\epsilon$ (2)
where $\mathcal{G}(\cdot)$ is a lightweight linear projection.
The Generalized Householder Update: To strictly implement the geometric capability of “erasing” the current state information, the update rule must include a subtractive term proportional to the current state. We formulate the final update as:
$\mathbf{x}_{l+1}=\mathbf{x}_{l}+\bm{\beta}\odot(\mathbf{v}_{mHC}-\alpha\cdot\text{Proj}_{\mathbf{x}}(\mathbf{x}_{l}))$ (3)
In our simplified implementation, we approximate the projection term $\text{Proj}_{\mathbf{x}}(\mathbf{x}_{l})$ with the normalized state itself, governed by a learnable scalar $\alpha$. This formulation allows:
  • Accumulation: When $\beta>0$ and $\mathbf{v}_{mHC}$ dominates.
  • Erasure/Reflection: When $\beta$ scales the negative term $\alpha\mathbf{x}_{l}$, effectively subtracting current features from the residual stream (Householder-like reflection).
@@ 3.5 Algorithm
[…]
10:4. Geometric Update (Householder)
@@ 4.2 Experiment 1: Rank Evolution Analysis
$\text{Rank}_{eff}(\mathbf{X}_{l})=\frac{\exp(H(\bm{\sigma}(\mathbf{X}_{l})))}{S}$ (4)
where $H(\bm{\sigma})=-\sum_{i}\hat{\sigma}_{i}\log\hat{\sigma}_{i}$ is the Shannon entropy of the normalized singular values $\hat{\sigma}_{i}=\sigma_{i}/\sum_{j}\sigma_{j}$.
@@ 4.4 Experiment 3: Beta Distribution Analysis
  • Early Layers: $\mathbb{E}[\beta]>0$ (accumulation mode for low-level features).
  • Deep Layers: Significant fraction of $\beta<0$ (erasure mode for semantic refinement).
  • Training Dynamics: The transition point should sharpen as training progresses.

### v3

@@ Abstract
Scaling Transformer architectures to extreme depth often leads to rank collapse: representations become redundant and degenerate despite modern normalization schemes. We argue this is fundamentally a geometric problem. Standard residual connections implicitly assume monotonic feature accumulation is beneficial, but provide no mechanism to constrain update directions or erase outdated information. As depth increases, this causes uncontrolled drift from the semantic manifold and representational collapse. We propose the Manifold-Geometric Transformer (MGT), a unified framework addressing these failures through two orthogonal principles. First, manifold-constrained hyper-connections (mHC) restrict residual updates to valid tangent space directions, preventing manifold drift. Second, deep delta learning (DDL) enables data-dependent, non-monotonic updates that support feature erasure rather than unconditional accumulation. Together, mHC controls update direction while DDL controls magnitude and sign, yielding stable geometric evolution across depth.
@@ 1 Introduction
Simultaneously, the additive nature implies a “write-only” memory mechanism. It is difficult for a network to learn to “erase” information (i.e., learn $\mathcal{F}(\mathbf{x})\approx-\mathbf{x}$) when context shifts. This results in the “residual accumulation” problem, where noise accumulates over depth.
@@ 2 Related Work
Rank Collapse in Deep Networks. Dong et al. (2021) provided seminal theoretical analysis showing that pure self-attention layers lose rank doubly exponentially with depth, fundamentally limiting representational capacity. This collapse occurs even with modern normalization techniques (Ba et al., 2016; Xiong et al., 2020), suggesting that architectural modifications beyond optimization tricks are necessary. Recent depth-scaling efforts (Wang et al., 2022; Touvron et al., 2021) have proposed specialized initialization and normalization schemes to train networks exceeding 100 layers, yet these approaches treat symptoms rather than addressing the geometric root cause of collapse.
Residual Connections and Depth. Since ResNets (He et al., 2016), residual connections have enabled training of deep networks by providing gradient highways. However, the standard formulation $\mathbf{x}_{l+1}=\mathbf{x}_{l}+\mathcal{F}(\mathbf{x}_{l})$ implicitly assumes additive updates in Euclidean space are universally beneficial. As Touvron et al. (2021) observed, this monotonic accumulation can lead to feature redundancy in very deep networks. Our work addresses this through explicit erasure mechanisms.
Manifold-Constrained Hyper-Connections (mHC). The recently proposed mHC framework (Xie et al., 2025) addresses stability in width-scaling by constraining hyper-connections to geometric manifolds (e.g., the Birkhoff polytope), preserving the identity mapping property of ResNets. We adopt this perspective as a “spatial regularizer” that projects updates onto the tangent space $T_{\mathbf{x}}\mathcal{M}$ of the data manifold. While Xie et al. (2025) focused on connection geometry, we argue that dynamics along these connections require complementary mechanisms to address feature redundancy.
Deep Delta Learning (DDL). Our dynamic update law builds on the recent Deep Delta Learning framework (Zhang et al., 2026), which generalizes residual connections via a learnable Delta Operator—a rank-1 perturbation of the identity matrix. The Delta Operator is formally defined as:
$\mathbf{A}(\mathbf{X})=\mathbf{I}-\beta(\mathbf{X})\mathbf{k}(\mathbf{X})\mathbf{k}(\mathbf{X})^{\top}$ (1)
where $\mathbf{k}(\mathbf{X})\in\mathbb{R}^{d}$ is a unit-normalized reflection direction vector and $\beta(\mathbf{X})\in[0,2]$ is a data-dependent gating scalar. The spectral analysis (Theorem 3.1 in Zhang et al. (2026)) reveals that $\mathbf{A}$ has (d1)(d-1) eigenvalues of $1$ spanning $\mathbf{k}^{\perp}$ and one eigenvalue of $(1-\beta)$ with eigenvector $\mathbf{k}$, yielding $\det(\mathbf{A})=1-\beta$. This enables three fundamental geometric regimes: identity preservation ($\beta\to 0$), orthogonal projection/forgetting ($\beta\to 1$, $\det\to 0$), and Householder reflection ($\beta\to 2$, $\det\to-1$). The complete Delta Residual Block combines this operator with a rank-1 value injection:
$\mathbf{X}_{l+1}=\mathbf{A}(\mathbf{X}_{l})\mathbf{X}_{l}+\beta(\mathbf{X}_{l})\mathbf{k}(\mathbf{X}_{l})\mathbf{v}(\mathbf{X}_{l})^{\top}=\mathbf{X}_{l}+\beta_{l}\mathbf{k}_{l}(\mathbf{v}_{l}^{\top}-\mathbf{k}_{l}^{\top}\mathbf{X}_{l})$ (2)
This additive form reveals synchronous “erase-and-write” semantics: the term $\mathbf{k}_{l}^{\top}\mathbf{X}_{l}$ represents old memory to be erased, while $\mathbf{v}_{l}^{\top}$ injects new information—both scaled by the same gate $\beta_{l}$. This recovers the classical Delta Rule from associative memory theory, reinterpreting network depth as iterative memory refinement.
@@ 3.1 Architecture Overview
We redesign the fundamental building block of the Transformer to explicitly separate feature generation from feature propagation. As illustrated in Figure 1, the proposed Manifold-Geometric Transformer (MGT) Block deviates from the standard Post-LayerNorm (Post-LN) structure by introducing a geometric processing stage before the residual addition.
@@
Figure 1: Architecture of the Manifold-Geometric Transformer (MGT) Block. The pipeline explicitly separates three phases: (1) Generation via LayerNorm and MHSA/FFN producing raw updates $\mathbf{V}_{raw}$, (2) Geometric Rectification via mHC projection $\Psi$ constraining updates to the tangent space, and (3) Dynamics via the DDL gate $\bm{\beta}(\mathbf{X}_{l})$ controlling update magnitude. The skip connection preserves identity mapping while the gated output $\mathbf{G}=\bm{\beta}\odot\mathbf{V}_{mHC}$ enables selective erasure and accumulation.
@@ 3.2 Preliminaries: The Manifold Hypothesis in Deep Layers
Let $\{\mathbf{X}_{l}\}_{l=0}^{L}$ denote the sequence of hidden states in a Transformer, where $\mathbf{X}_{l}\in\mathbb{R}^{S\times D}$ represents $S$ tokens with $D$-dimensional features. We posit that valid semantic representations do not populate the entire Euclidean space $\mathbb{R}^{D}$, but rather lie on a lower-dimensional Riemannian manifold $\mathcal{M}\subset\mathbb{R}^{D}$. A naive step $\mathbf{X}_{l}+\mathbf{V}_{l}$ often forces the representation off the manifold ($\mathbf{X}_{l+1}\notin\mathcal{M}$). To maintain structural integrity, the update vector must be constrained to the local tangent space $T_{\mathbf{X}_{l}}\mathcal{M}$.
3.3 The Delta Operator
The key innovation of Deep Delta Learning (Zhang et al., 2026) is to generalize the classical Householder reflection (see Appendix A) by replacing the constant factor $2$ with a learnable, data-dependent scalar gate $\beta(\mathbf{X})$.
Definition 3.1 (Delta Operator).
For a unit vector $\mathbf{k}\in\mathbb{R}^{d}$ and scalar $\beta\in\mathbb{R}$, the Delta Operator is:
$\mathbf{A}(\beta,\mathbf{k})=\mathbf{I}-\beta\mathbf{k}\mathbf{k}^{\top}$ (3)
This constitutes a rank-1 perturbation of the identity. When $\beta=2$, it recovers the Householder reflector.
The geometric behavior of the Delta Operator is characterized by its spectrum:
Theorem 3.2 (Spectral Decomposition).
The spectrum of $\mathbf{A}=\mathbf{I}-\beta\mathbf{k}\mathbf{k}^{\top}$ is $\sigma(\mathbf{A})=\{1,\ldots,1,1-\beta\}$ with (d1)(d-1) eigenvalues of $1$ in $\mathbf{k}^{\perp}$ and eigenvalue $(1-\beta)$ for eigenvector $\mathbf{k}$. Consequently, $\det(\mathbf{A})=1-\beta$.
This spectral structure yields three fundamental geometric regimes as $\beta$ varies in $[0,2]$:
Proposition 3.3 (Geometric Regimes).
(i) Identity ($\beta\to 0$): $\mathbf{A}\to\mathbf{I}$, preserving input and enabling layer skipping. (ii) Projection ($\beta\to 1$): $\mathbf{A}\to\mathbf{P}_{\mathbf{k}^{\perp}}$, erasing the $\mathbf{k}$-component ($\det\to 0$). (iii) Reflection ($\beta\to 2$): $\mathbf{A}\to\mathbf{H}_{\mathbf{k}}$, inverting the $\mathbf{k}$-component ($\det\to-1$).
The complete Delta Residual Block combines this operator with value injection:
$\mathbf{X}_{l+1}=\mathbf{A}(\mathbf{X}_{l})\mathbf{X}_{l}+\beta_{l}\mathbf{k}_{l}\mathbf{v}_{l}^{\top}=\mathbf{X}_{l}+\beta_{l}\mathbf{k}_{l}(\mathbf{v}_{l}^{\top}-\mathbf{k}_{l}^{\top}\mathbf{X}_{l})$ (4)
The additive form reveals “erase-and-write” semantics: $\mathbf{k}_{l}^{\top}\mathbf{X}_{l}$ is erased while $\mathbf{v}_{l}^{\top}$ is written, both scaled by $\beta_{l}$. This recovers the classical Delta Rule from associative memory (proofs in Appendix A).
@@ 3.4 Manifold-Constrained Hyper-Connections (mHC)
The goal of mHC is to rectify the raw update $\mathbf{V}_{raw}$ generated by the sub-layers (MHSA or FFN). Since computing the exact tangent space via eigendecomposition is computationally prohibitive, we employ an efficient Soft Subspace Approximation.
@@
$\mathbf{V}_{mHC}=\mathbf{V}_{raw}\odot\sigma(\text{LN}(\mathbf{X}_{l}\mathbf{W}_{gate}))$ (5)
where $\mathbf{W}_{gate}\in\mathbb{R}^{D\times D}$ projects the input to semantic importance scores. While not a strict orthogonal projection in the linear algebra sense, this acts as a soft manifold constraint, suppressing feature dimensions (noise subspace) that deviate from the current semantic trajectory defined by $\mathbf{X}_{l}$.
3.5 MGT Integration: Combining mHC with DDL
We now describe how MGT integrates DDL with mHC to achieve stable geometric updates.
3.5.1 The MGT Delta Controller
In MGT, we treat the mHC-rectified output $\mathbf{V}_{mHC}$ as the value branch in the Delta Residual Block (Eq. 4). We introduce a Delta Controller $\bm{\beta}(\mathbf{X}_{l})$:
$\bm{\beta}=\lambda\cdot\tanh(\mathbf{X}_{l}\mathbf{W}_{\beta}+\mathbf{b}_{\beta})+\epsilon$ (6)
where $\mathbf{W}_{\beta}\in\mathbb{R}^{D\times D}$ and $\mathbf{b}_{\beta}\in\mathbb{R}^{D}$ are learnable parameters, $\lambda$ scales the output range, and $\epsilon$ provides numerical stability.
Remark 3.4 (Range Extension).
The original DDL uses $\beta\in[0,2]$ via sigmoid ($\beta=2\sigma(\cdot)$). Our $\tanh$ parameterization with offset $\epsilon$ yields $\beta\in[-\lambda+\epsilon,\lambda+\epsilon]$, extending the geometric capability to include negative gates for more aggressive signed updates. Setting $\epsilon=0$ and $\lambda=1$ recovers the standard range [1,1][-1,1].
3.5.2 The MGT Update Rule
Following the additive form (Eq. 4), the MGT update implements synchronous erase-and-write semantics:
$\mathbf{X}_{l+1}=\mathbf{X}_{l}+\bm{\beta}\odot(\mathbf{V}_{mHC}-\alpha\cdot\mathbf{X}_{l})$ (7)
where $\alpha\in\mathbb{R}$ is a learnable scalar controlling the erasure strength. In the full DDL formulation, this corresponds to the projection $\text{Proj}_{\mathbf{k}}(\mathbf{X}_{l})$; we simplify by approximating with the scaled state $\alpha\cdot\mathbf{X}_{l}$.
3.5.3 Synergy Analysis
The combination of mHC and DDL provides complementary geometric controls:
Proposition 3.5 (mHC-DDL Synergy).
Let $\mathbf{V}_{mHC}=\Psi(\mathbf{V}_{raw})$ be the mHC-rectified update. The MGT update rule provides:
  1. 1.
    Direction Control (mHC): Ensures $\mathbf{V}_{mHC}$ lies approximately in the tangent space $T_{\mathbf{X}}\mathcal{M}$, preventing manifold drift.
  2. 2.
    Magnitude Control (DDL): The gate $\beta$ modulates the step size along the rectified direction. When $\mathbf{V}_{mHC}>\alpha\mathbf{X}_{l}$ (value-dominant), the update accumulates new features; when $\mathbf{V}_{mHC}<\alpha\mathbf{X}_{l}$ (projection-dominant), the update erases existing features. Negative $\beta$ inverts the update direction entirely.
  3. 3.
    Gradient Preservation: By Theorem 3.2, (d1)(d-1) eigenvalues remain exactly $1$, ensuring unimpeded gradient flow through $\mathbf{k}^{\perp}$ even under aggressive erasure.
This orthogonal decomposition—mHC controlling where to update and DDL controlling how much—enables stable “Erasure-and-Write” dynamics essential for ultra-deep scaling.
@@ 3.6 Algorithm
9:4. Geometric Update (Householder)
@@ 4.2 Experiment 1: Rank Evolution Analysis
$\text{Rank}_{eff}(\mathbf{X}_{l})=\frac{\exp(H(\bm{\sigma}(\mathbf{X}_{l})))}{\min(S,D)}$ (8)
where $H(\bm{\sigma})=-\sum_{i}\hat{\sigma}_{i}\log\hat{\sigma}_{i}$ is the Shannon entropy of the normalized singular values $\hat{\sigma}_{i}=\sigma_{i}/\sum_{j}\sigma_{j}$. The normalization by $\min(S,D)$ ensures $\text{Rank}_{eff}\in[0,1]$.
@@ 4.4 Experiment 3: Beta Distribution Analysis
  • Early Layers: $\mathbb{E}[\beta]>0$ (standard accumulation mode for low-level feature extraction).
  • Deep Layers: Significant fraction of $\beta<0$ (update reversal mode), indicating the model learns to “undo” or refine early-layer representations. Note: in MGT, erasure also occurs when $\mathbf{V}_{mHC}<\alpha\mathbf{X}_{l}$ regardless of $\beta$ sign.
  • Training Dynamics: The transition point (layer at which $\mathbb{E}[\beta]$ changes sign) should sharpen as training progresses.
@@ 5 Conclusion
Appendix A Mathematical Proofs and Derivations
This appendix provides complete mathematical foundations and proofs for the results stated in the main text.
A.1 The Householder Transformation
Definition A.1 (Householder Matrix).
For a non-zero vector $\mathbf{k}\in\mathbb{R}^{d}$, the Householder matrix is:
$\mathbf{H}_{\mathbf{k}}=\mathbf{I}-2\frac{\mathbf{k}\mathbf{k}^{\top}}{\|\mathbf{k}\|_{2}^{2}}$ (10)
For unit vectors ($\|\mathbf{k}\|_{2}=1$), this simplifies to $\mathbf{H}_{\mathbf{k}}=\mathbf{I}-2\mathbf{k}\mathbf{k}^{\top}$.
Lemma A.2 (Properties of Householder Matrices).
For any Householder matrix $\mathbf{H}_{\mathbf{k}}$: (1) Symmetry: $\mathbf{H}_{\mathbf{k}}=\mathbf{H}_{\mathbf{k}}^{\top}$; (2) Orthogonality: $\mathbf{H}_{\mathbf{k}}^{\top}\mathbf{H}_{\mathbf{k}}=\mathbf{I}$; (3) Involution: $\mathbf{H}_{\mathbf{k}}^{2}=\mathbf{I}$; (4) Determinant: $\det(\mathbf{H}_{\mathbf{k}})=-1$.
Proof.
Properties (1)-(3) follow directly from the definition. For (4), $\mathbf{H}_{\mathbf{k}}$ has eigenvalue 1-1 with eigenvector $\mathbf{k}$ and eigenvalue $1$ with multiplicity (d1)(d-1) in $\mathbf{k}^{\perp}$, hence $\det(\mathbf{H}_{\mathbf{k}})=(-1)\cdot 1^{d-1}=-1$. ∎
A.2 Proof of Theorem 3.2 (Spectral Decomposition)
Proof.
We prove that $\mathbf{A}=\mathbf{I}-\beta\mathbf{k}\mathbf{k}^{\top}$ has spectrum $\sigma(\mathbf{A})=\{1,\ldots,1,1-\beta\}$.
Part 1: For any $\mathbf{u}\in\mathbf{k}^{\perp}$ (i.e., $\mathbf{k}^{\top}\mathbf{u}=0$):
$\mathbf{A}\mathbf{u}=(\mathbf{I}-\beta\mathbf{k}\mathbf{k}^{\top})\mathbf{u}=\mathbf{u}-\beta\mathbf{k}(\mathbf{k}^{\top}\mathbf{u})=\mathbf{u}$ (11)
Thus every vector in $\mathbf{k}^{\perp}$ is an eigenvector with eigenvalue $1$. Since $\dim(\mathbf{k}^{\perp})=d-1$, this accounts for (d1)(d-1) eigenvalues of $1$.
Part 2: For the vector $\mathbf{k}$ itself (with $\|\mathbf{k}\|_{2}=1$):
$\mathbf{A}\mathbf{k}=(\mathbf{I}-\beta\mathbf{k}\mathbf{k}^{\top})\mathbf{k}=\mathbf{k}-\beta\mathbf{k}(\mathbf{k}^{\top}\mathbf{k})=(1-\beta)\mathbf{k}$ (12)
Thus $\mathbf{k}$ is an eigenvector with eigenvalue $(1-\beta)$.
Since we have identified $d$ linearly independent eigenvectors, the spectral decomposition is complete. The determinant formula $\det(\mathbf{A})=1^{d-1}\cdot(1-\beta)=1-\beta$ follows immediately. ∎
A.3 Additional Corollaries
Corollary A.3 (Orthogonality Condition).
The Delta Operator $\mathbf{A}$ is orthogonal iff $|1-\beta|=1$, i.e., $\beta\in\{0,2\}$.
Proof.
A matrix is orthogonal iff all eigenvalues have absolute value $1$. Since (d1)(d-1) eigenvalues are exactly $1$, we require $|1-\beta|=1$, yielding $\beta=0$ or $\beta=2$. ∎
A.4 Proof of Additive Form (Eq. 4)
Proof.
Expanding $\mathbf{X}_{l+1}=\mathbf{A}(\mathbf{X}_{l})\mathbf{X}_{l}+\beta_{l}\mathbf{k}_{l}\mathbf{v}_{l}^{\top}$:
$\mathbf{X}_{l+1}$ $=(\mathbf{I}-\beta_{l}\mathbf{k}_{l}\mathbf{k}_{l}^{\top})\mathbf{X}_{l}+\beta_{l}\mathbf{k}_{l}\mathbf{v}_{l}^{\top}$ (13)
$=\mathbf{X}_{l}-\beta_{l}\mathbf{k}_{l}\mathbf{k}_{l}^{\top}\mathbf{X}_{l}+\beta_{l}\mathbf{k}_{l}\mathbf{v}_{l}^{\top}$ (14)
$=\mathbf{X}_{l}+\beta_{l}\mathbf{k}_{l}(\mathbf{v}_{l}^{\top}-\mathbf{k}_{l}^{\top}\mathbf{X}_{l})$ (15)
This reveals the “erase-and-write” semantics of the Delta Rule. ∎
A.5 Vector Case
When $d_{v}=1$ (vector hidden states $\mathbf{x}\in\mathbb{R}^{d}$), the update simplifies to:
$\mathbf{x}_{l+1}=\mathbf{x}_{l}+\beta_{l}(v_{l}-\mathbf{k}_{l}^{\top}\mathbf{x}_{l})\mathbf{k}_{l}$ (16)
where $v_{l}\in\mathbb{R}$ is a scalar and $(v_{l}-\mathbf{k}_{l}^{\top}\mathbf{x}_{l})$ acts as a data-dependent step size.
A.6 Dimensional Correspondence for Transformers
In the Transformer context with hidden states $\mathbf{X}_{l}\in\mathbb{R}^{S\times D}$, the Delta Operator acts on the hidden dimension $D$. We identify $d=D$ (hidden dimension) and $d_{v}=1$ for the simplified vector case, or apply the operator token-wise with shared parameters.